OpenStudy (anonymous):

what type of asymptote would g(x)=1/3* 4^x have and where? Also given f(x)= 4^x?

OpenStudy (aum):

What happens to g(x) as x approaches -infinity?

OpenStudy (anonymous):

I'm looking at the graph right now but I'm not sure haha

OpenStudy (anonymous):

I don't understand at all how to find the retricemptote or where to see it. I'm not sure if I should be able to tell by looking at a graph or solving an equation

OpenStudy (aum):

I have heard the term asymptote. How do they define retricemptote in your book/notes?

OpenStudy (anonymous):

oh my gosh sorry that was a typo haha

OpenStudy (anonymous):

asymptote

OpenStudy (anonymous):

wow sorry about that

OpenStudy (aum):



OpenStudy (anonymous):

Yes I got that graph also

OpenStudy (aum):

As x approaches negative infinity, g(x) approaches zero. It never quite becomes zero but gets closer and closer to zero which is the x axis or y = 0. So the x axis or the line y = 0 is a horizontal asymptote. There are no vertical asymptotes for this graph.

OpenStudy (anonymous):

so it's zero because it is almost 0? Than how would you do it for 4^(x-5) +19? and can you only tell whether it's vertical or horizontal just by looking at a graph or can you do it by looking at a equation too

OpenStudy (anonymous):

How do you give medals because I'll give you one lol

OpenStudy (aum):

You can find many asymptotes by looking at the graph. But the proper way to do it is by finding the limit of f(x) as x approaches \(\pm \infty\). You find the vertical asymptotes by looking at the denominator and see what vale of x will make the denominator zero.

OpenStudy (anonymous):

So if there is no denomintor..

OpenStudy (anonymous):

Can you also help me find the range of this graph also?

OpenStudy (aum):

Some of the examples here might help: http://www.purplemath.com/modules/asymtote.htm Or you can type asymptotes in YouTube and check out a couple of vids.

OpenStudy (aum):

g(x)=1/3* 4^x never goes below the x-axis and so it is never negative. It approaches zero but never gets to zero. So the lowest value it can have is close to zero. You indicate this by an open parenthesis: (0, g(x) approaches +infinity as x approaches +infinity. So the range of g*x) is: \(\large (0, \infty)\).

OpenStudy (anonymous):

(0,inf) is wrong for the range I tried that already :(

OpenStudy (aum):

Is this the correct function? \(\large g(x) = \frac 13 4^x\)

OpenStudy (anonymous):

4^(x-5) +19

OpenStudy (anonymous):

And I don't understand how to find vertical asymptotes for non rational functions, where there is no denominator